A constant net force of 200 N is applied to a cart in frictionless surface from rest to a velocity of 4m/s in s. What is the mass of the cart!

1
Expert's answer
2021-11-30T09:49:41-0500

Let's first find the acceleration of the cart. Unfortunately, we don't know how long the cart accelerates by the force. Let's suppose that it takes 10 s to reach the velocity of 4 m/s. Then, we get:


"a=\dfrac{v-v_0}{t}=\dfrac{4\ \dfrac{m}{s}-0}{10\ s}=0.4\ \dfrac{m}{s^2}."

Finally, we can find the mass of the cart from the Newton's Second Law of Motion:


"F=ma,""m=\dfrac{F}{a}=\dfrac{200\ N}{0.4\ \dfrac{m}{s^2}}=500\ kg."